OpenStudy (anonymous):

A swimmer wants to cross 500 m wide river flowing at 5km/h. his speed with respect to water is 3km/h. the man has to reach the point directly opposite to his starting point. If he reaches the other shore somewhere else, he has to walk to the point . Find the minimum distance that he has to walk. (swimmer makes an angle thita)

OpenStudy (anonymous):

The swimmer makes angle thita with the river flow .Therefore component of velocity in horizontal dir.=5+3cos thita component of velocity in vertical dir.=3sin thita Therefore,time to cross river =0.5/3sin thita=1/6sin thita min. distance swimmer has to walk =(5+3cos thita)(1/6sin thita)=H..............(1) for min. distance dH/d thita =0 why?? ------------- dH/d thita =0 ??

OpenStudy (anonymous):